OpenStudy (anonymous):

Find the limit of sin(x^2)/x as x approaches to 0.

OpenStudy (isaiah.feynman):

Use L'hospital's rule. Take the derivatives of the numerator and denominator SEPARATELY.

OpenStudy (anonymous):

I did. But I got 2x(cos x^2)/1, what's next?

OpenStudy (isaiah.feynman):

Substitute x=0 into the equation.

OpenStudy (anonymous):

I got 0/1=0, so is that the answer?

OpenStudy (anonymous):

its simple sinx^2/x=(sinx.sinx)/x =(sinx/x)sinx applying limit as x tends to 0, we know sinx/x=1 thus =1*sinx since x is 0 thus sinx=0 and 1*0=0 hence answer is =1*0=0